# Nesterov’s Accelerated Jacobi-Type Methods for Large-scale Symmetric Positive Semidefinite Linear Systems

### v1

@@ 2.2 An adaptive restarting strategy
The template of our restarted algorithm is described in Algorithm 2. Here, we briefly explain the ideas of the proposed restarting scheme. Observe that the updating rule of $y^{t+1}$ reads $y^{t+1}=x^{t}+\frac{\alpha_{t}-1}{\alpha_{t+1}}(x^{t}-x^{t-1})$, which amounts to adding the momentum $\frac{\alpha_{t}-1}{\alpha_{t+1}}(x^{t}-x^{t-1})$ to $x^{t}$. The significance of this momentum $\frac{\alpha_{t}-1}{\alpha_{t+1}}\in[0,1]$ approaches 1 quickly due to the fact that $\alpha_{t}\geq\frac{t+1}{2}$, for $t\geq 0$. This can cause the generated iterates $x^{t}$ and $y^{t}$ to oscillate (i.e., overshoot) near the optimal solution $x^{*}$, which further slows down the convergence speed of the algorithm, based on our empirical observation,. To address this concern, an adaptive restarting strategy that intermittently resets the momentum’s significance back to zero, utilizing the point generated from the preceding iteration as the new starting point (by setting $\alpha_{t+1}=1$, $\alpha_{t}=0$, $y^{t+1}=x^{t-1}$, and $x^{t}=x^{t-1}$). There remains one critical issue to be addressed: what is the right time to perform restarting? There is no affirmative answer to this question, and the approach we employed is somewhat heuristic in nature. On the one hand, when the monitored condition $\left\langle Qy^{t}-b,x^{t}-x^{t-1}\right\rangle\geq 0$ is satisfied, we see that the angle between $\nabla f(y^{t})$ and the direction of motion $x^{t}-x^{t-1}$ is acute. Intuitively, $x^{t}$ seems to make the objective function increase. So this condition can be a heuristic signal for us to trigger a restart. On the other hand, we do not want to perform the restarting too frequently in order to allow the algorithm to make significant progress. Moreover, oscillation may only be moderate, and a restart could destroy the fast convergence speed. Hence, similar to [18], we set a prohibition period and allow the algorithm to restart only when the number of iterations between the $(\ell-1)$-th restart and the $\ell$-th restart is larger than $K_{\ell}$, where $K_{\ell}$ is updated as $K_{\ell+1}=2K_{\ell}$ for $\ell\geq 0$ and $K_{0}\geq 2$ is specified by the user.
@@
First, assume that the restart does not occur until the $(t+1)$-th iteration. From Lemma 2.1, it is easy to check that
[…]
In particular, for $j=1$, we have
@@
Summing over $j=1,2,\dots,t-1$, we obtain
@@
Recall that $\alpha_{1}=1$ and that
@@
[…]
Thus, the statements of the lemma follows easily and hence the proof is completed.
The above lemma indicates that the sequence $\{f(x^{\ell,0})\}$ is nonincreasing. Using this fact and assuming that $Q\in\mathbb{S}_{++}^{n}$, we are able to establish the convergence of Algorithm 2.
@@
Let the sequence $\{x^{t}:=x^{\ell,k_{\ell}}\}$ be generated by Algorithm 2. Suppose that $Q$ is SPD. Then, it holds that
@@
Since $Q\in\mathbb{S}_{++}^{n}$, we see that $f$ is coercive. Furthermore, the level set
$\mathcal{L}(x^{0}):=\{x\in\mathbb{R}^{n}\;:\;f(x)\leq f(x^{0})\}$
is bounded and there exist a constant $R>0$ such that
$R\geq\sup_{x\in\mathcal{L}(x^{0})}\left\lVert x-x^{*}\right\rVert_{S}.$
@@
[…]
@@
We emphasize that the positive definiteness of $Q$ is necessary in our analysis since it ensures the existence of the positive constant $R$ which serves as an upper bound for the diameter of the level set $\mathcal{L}(x^{0}):=\{x\in\mathbb{R}^{n}\;:\;f(x)\leq f(x^{0})\}$. In practice, we observe that Algorithm 2 still works efficiently for linear systems with $Q\in\mathbb{S}_{+}^{n}$. Moreover, the convergence rate given by Theorem 2.9 introduces a $\log_{2}^{2}(t)$ term which is worse than the one given by Theorem 2.5. Although Algorithm 2 has a poorer worst-case convergence rate than that of Algorithm 1, it has better practical numerical performance empirically. It remains an open problem whether the restarting scheme admits better theoretical convergence properties.
@@ 3 A new proximal term for parallelization
Now, we consider choosing a matrix $J$ as a block-diagonal matrix, i.e., $J=\operatorname{diag}(J^{1,1},\dots,J^{s,s})$ with $J^{i,i}$ being a symmetric positive definite matrix. Let $S:=J-Q$. Then, to get $x^{t}$ in Line 3 of Algorithm 2, it suffices to solve the following linear system
@@
which is indeed a Jacobi-type updating rule and can readily be implemented in a parallel manner. The template of the parallel version using the above matrix $S$ with an adaptive restarting strategy is given in Algorithm 3.
1:  Input: Initial points $x^{0}=y^{1}\in\mathbb{R}^{n}$, $\alpha_{0}=0,\;\alpha_{1}=1$, block-diagonal matrix $J:=\operatorname{diag}(J^{1,1},\dots,J^{s,s})\in\mathbb{S}_{++}^{n}$, $\textrm{IsRestart}\in\{\textrm{True},\textrm{False}\}$, $\ell=0$, $K_{\ell}\geq 2$ and $K_{\textrm{re}}=0$.
@@
to get a block diagonal $J$.
@@ 4 Numerical experiments
In this section, we conduct numerical experiments to verify the efficiency and scalability of the proposed algorithm, i.e., Algorithm 3 with an adaptive restarting strategy. The outline and main goals of our numerical experiments are summarized as follows:
@@
    We first construct a set of simple strictly diagonal dominant SPD systems for which the classical Jacobi and weighted Jacobi methods are both guaranteed to be convergent. The purpose of this set of numerical tests is to show that our proposed algorithm outperforms the classical Jacobi-type methods.
@@
    We repeat the above tests on some SPD systems with real data sets in order to evaluate the performance of the proposed algorithm on SPD systems.
@@
In our experiment, we only test Algorithm 3 with $J$ chosen as in (7), since, based on our numerical experience, Algorithm 3 with $J$ chosen as in (8) shares a similar performance as the former one. We terminate all the tested methods based on the value of the relative residue norm, i.e., $\frac{\left\lVert b-Qx^{t}\right\rVert}{\left\lVert b\right\rVert}$, where $x^{t}$ is the generated iterate at $t$-th iteration of an algorithm. In particular, we set the termination tolerance tol to $10^{-4}$. We also set the maximum number of iterations allowed, maxiter, to be $5000$.
@@ 4.1 Strictly diagonal dominant systems
Figure 2: Computational results for strictly diagonally dominant systems with $\texttt{maxiter}=5000$ and $\texttt{tol}=10^{-4}$.
From Figure 2, we observe that both the Jacobi method and the weighted Jacobi method, denoted as jacobi and w-jacobi, respectively, are convergent and admit linear convergence rates. Moreover, the weighted Jacobi method with the optimal weight has a faster convergence rate than that of the classical Jacobi method. However, both methods are not able to solve most problems to the desired accuracy. On the other hand, the proposed algorithm, denoted as acc-jacobi, significantly outperforms the other two methods and is able to solve all the problems successfully to the desired accuracy. This provides strong evidence that the Nesterov’s accelerated technique equipped with the adaptive restarting starting strategy is indeed efficient.
@@ 4.2 Laplacian systems
$[L(G)]_{ij}:=\left\{\begin{array}[]{ll}\textrm{deg}(v_{i}),&\textrm{if $i=j$},\\ -1,&\textrm{if $i\neq j$ and $(i,j)\in E$},\\ 0,&\textrm{otherwise,}\end{array}\right.\quad 1\leq i,j\leq n,$
where $\textrm{deg}(v_{i})$ denotes the degree of the node $v_{i}$, for $1\leq i\leq n$. It is well-known that the Laplacian matrix $L(G)$ is symmetric positive semidefinite and singular with $\textrm{rank}(L(G))=n-1$. Solving Laplacian linear systems is a fundamental computational task with fruitful applications in a wide range of fields. The classical Jacobi method and the weighted Jacobi method are not guaranteed to be convergent when solving Laplacian systems. In fact, our numerical tests show that both methods are divergent empirically. Therefore, we only compare the proposed algorithm (i.e., acc-jacobi) with the CG method and the PCG method with a simple diagonal preconditioner. We select a set of graphs arising from real-world applications in the SuiteSparse Matrix Collection  [^2] Available at https://sparse.tamu.edu/. with various sizes. The corresponding numerical results are presented in Figure 3.
@@
Figure 3: Computational results for Laplacian systems with $\texttt{maxiter}=5000$ and $\texttt{tol}=10^{-4}$.
We can observe from Figure 3 that all the three methods are quite efficient and are able to solve all the tested problems to the specified accuracy. We can also see that acc-jacobi and pcg have shown superior performance when compared to the pure CG method, i.e., cg. However, acc-jacobi performs slightly worse than the pcg method. In particular, the number of iterations taken by acc-jacobi is sometimes almost twice the number of iterations taken by pcg.
4.3 SPD systems
We also test acc-jacobi, cg and pcg on SPD systems whose coefficient matrices are also collected from the SuiteSparse Matrix Collection. These matrices are real-world data coming from different practical science and engineering problems. The numerical results presented in Figure 4 and Figure 5 show that acc-jacobi and pcg have comparable performance, and they generally perform better than cg.
(a) bcsstk11
(b) bcsstk12
(c) bcsstk13
(d) bcsstk14
(e) bcsstk24
(f) bcsstk27
(g) bcsstk28
(h) ex13
(i) msc01050
(j) msc01440
(k) msc10848
(l) nasa1824
Figure 4: Computational results for PSD systems collected from the SuiteSparse collection with $\texttt{maxiter}=5000$ and $\texttt{tol}=10^{-4}$.
(a) nasa1824
(b) nasa2910
(c) nd3k
(d) nd6k
(e) nd12k
(f) Trefethen-2000
Figure 5: (Continued) Computational results for systems from the SuiteSparse collection with $\texttt{maxiter}=5000$ and $\texttt{tol}=10^{-4}$.

### v2

@@ 2.2 An adaptive restarting strategy
The template of our restarted algorithm is described in Algorithm 2. Here, we briefly explain the ideas of the proposed restarting scheme. Observe that the updating rule of $y^{t+1}$ reads $y^{t+1}=x^{t}+\frac{\alpha_{t}-1}{\alpha_{t+1}}(x^{t}-x^{t-1})$, which amounts to adding the momentum $\frac{\alpha_{t}-1}{\alpha_{t+1}}(x^{t}-x^{t-1})$ to $x^{t}$. The significance of this momentum $\frac{\alpha_{t}-1}{\alpha_{t+1}}\in[0,1]$ approaches 1 quickly due to the fact that $\alpha_{t}\geq\frac{t+1}{2}$, for $t\geq 0$. This can cause the generated iterates $x^{t}$ and $y^{t}$ to oscillate (i.e., overshoot) near the optimal solution $x^{*}$, which further slows down the convergence speed of the algorithm, based on our empirical observation. To address this concern, an adaptive restarting strategy that intermittently resets the momentum’s significance back to zero, utilizing the point generated from the preceding iteration as the new starting point (by setting $\alpha_{t+1}=1$, $\alpha_{t}=0$, $y^{t+1}=x^{t-1}$, and $x^{t}=x^{t-1}$). There remains one critical issue to be addressed:
What is the right time to perform restarting?
There is no affirmative answer to this question, and the approach we employed is somewhat heuristic in nature. On the one hand, when the monitored condition in Line 4 of Algorithm 2, i.e., $\left\langle Qy^{t}-b,x^{t}-x^{t-1}\right\rangle\geq 0$ is satisfied, we see that the angle between $\nabla f(y^{t})$ and the direction of motion $x^{t}-x^{t-1}$ is acute. Intuitively, $x^{t}$ seems to make the objective function increase. So this condition can be a heuristic signal for us to trigger a restart. On the other hand, we do not want to perform the restarting too frequently in order to allow the algorithm to make significant progress. Moreover, oscillation may only be moderate, and a restart could destroy the fast convergence speed. Hence, similar to [20], we set a prohibition period and allow the algorithm to restart only when the number of iterations between the $(\ell-1)$-th restart and the $\ell$-th restart is larger than $K_{\ell}$, where $K_{\ell}$ is updated as $K_{\ell+1}=2K_{\ell}$ for $\ell\geq 0$ and $K_{0}\geq 2$ is specified by the user.
@@
First, we assume that the restart does not occur until the $(t+1)$-th iteration. From Lemma 2.1 and using the fact that
$\left\langle\mathbf{b}-\mathbf{a},S(\mathbf{b}-\mathbf{a})\right\rangle+2\left\langle\mathbf{a}-\mathbf{c},S(\mathbf{b}-\mathbf{a})\right\rangle=\left\langle\mathbf{b}-\mathbf{c},S(\mathbf{b}-\mathbf{c})\right\rangle-\left\langle\mathbf{a}-\mathbf{c},S(\mathbf{a}-\mathbf{c})\right\rangle$
for all $\mathbf{a},\mathbf{b},\mathbf{c}\in\mathbb{R}^{n}$, it is easy to check that
$f(x^{j})-f(x)\leq\frac{1}{2}\left\langle x-y^{j},S(x-y^{j})\right\rangle-\frac{1}{2}\left\langle x-x^{j},S(x-x^{j})\right\rangle,\quad\forall\;0\leq j\leq t,$
by letting $\mathbf{a}=y^{t}$, $\mathbf{b}=x^{j}$, and $\mathbf{c}=x$. In particular, for $j=1$ and $x=x^{0}$, we have
@@
Summing the above over $j=1,2,\dots,t-1$, we obtain that
@@
Recall that $\alpha_{1}=1$, we see that the second term on the right-hand-side vanishes. Moreover, one can observe from the relation $\alpha_{t+1}=(1+\sqrt{1+4\alpha_{t}^{2}})/2$ that
@@
This shows that the objective values are always not larger than the objective value at the initial piont if the restart is not performed.
@@
$f(x^{\ell,k_{\ell}})\leq f(x^{\ell,0}),\quad\forall\;0\leq k_{\ell}\leq\widetilde{K}_{\ell}$.
In particular, when $k_{\ell}=\widetilde{K}_{\ell}$, it holds that
$f(x^{\ell+1,0})=f(x^{\ell,\widetilde{K}_{\ell}})\leq f(x^{\ell,0}),\quad\forall\ell\,\geq 0.$
Thus, the proof is completed.
The above lemma indicates that the sequence $\{f(x^{\ell,0})\}$ is nonincreasing. Using this fact and assuming that $Q\in\mathbb{S}_{++}^{n}$, we are able to establish the convergence of Algorithm 2. In this case, one can see that the function $f$ is coercive and the level set
$\mathcal{L}(x^{0}):=\{x\in\mathbb{R}^{n}\;:\;f(x)\leq f(x^{0})\}$
is bounded. Consequently, there exist a constant $R>0$ such that
$R\geq\sup_{x\in\mathcal{L}(x^{0})}\left\lVert x-x^{*}\right\rVert_{S}.$
@@
Let the sequence $\{x^{t}:=x^{\ell,k_{\ell}}\}$ be generated by Algorithm 2. Suppose that $Q\in\mathbb{S}_{++}^{n}$. Then, it holds that
@@
$f(x^{\ell,k_{\ell}})-f(x^{*})\leq\frac{2\left\lVert x^{\ell,0}-x^{*}\right\rVert_{S}^{2}}{(k_{\ell}+1)^{2}}\leq\frac{2R^{2}}{(k_{\ell}+1)^{2}},\quad\forall\;\ell\geq 0,\;0\leq k_{\ell}\leq\widetilde{K}_{\ell}$,
@@
We emphasize that the positive definiteness of $Q$ is necessary in our analysis since it ensures the existence of the positive constant $R$ which serves as an upper bound for the diameter of the level set $\mathcal{L}(x^{0}):=\{x\in\mathbb{R}^{n}\;:\;f(x)\leq f(x^{0})\}$. In practice, we observe that Algorithm 2 still works efficiently for linear systems with $Q\in\mathbb{S}_{+}^{n}$. We should emphasize that the condition $\left\langle Qy^{t}-b,x^{t}-x^{t-1}\right\rangle\geq 0$ is not used in the convergence analysis. Instead, it is monitored in the algorithm as a heuristic to decide when to restart the extrapolation, which we found to improve practical efficiency. In contrast, the key to our convergence proof is the ability to avoid frequent restarts, which is ensured by enforcing the condition $t>K+K_{\ell}$. Moreover, the convergence rate given by Theorem 2.9 introduces a $\log_{2}^{2}(t)$ term which is worse than the one given by Theorem 2.5. Although Algorithm 2 has a poorer worst-case convergence rate than that of Algorithm 1, it has better practical numerical performance empirically. It remains an open problem whether the restarting scheme admits better theoretical convergence properties. Moreover, for a given problem, determining the optimal timing for restarting remains an open question that warrants further investigation. More advanced analytical tools and fully exploring the problem structure are required to bridge these gaps, which we leave as a direction for future research.
@@ 3 A new proximal term for parallelization
Now, we consider choosing a matrix $J$ as a block-diagonal matrix, i.e., $J=\operatorname{diag}(J^{1,1},\dots,J^{s,s})$ with $J^{i,i}$ being a symmetric positive definite matrix. Let S:=JQS:=J-Q. Then, to get $x^{t}$ in Line 3 of Algorithm 2, it suffices to solve the following linear system
@@
which is indeed a Jacobi-type updating rule and can readily be implemented in a parallel manner. Assuming that $S$ is positive semidefinite, the template of the parallel version using the above matrix $S$ with an adaptive restarting strategy can be given in Algorithm 3.
1:Input: Initial points $x^{0}=y^{1}\in\mathbb{R}^{n}$, $\alpha_{0}=0,\;\alpha_{1}=1$, the matrix $J:=\operatorname{diag}(J_{1},\dots,J_{s})\in\mathbb{S}_{++}^{n}$ defined by eq.(7) or (8), $\textrm{IsRestart}\in\{\textrm{True},\textrm{False}\}$, $\ell=0$, $K_{\ell}\geq 2$ and $K_{\textrm{re}}=0$.
@@
to get a block diagonal $J$. For both cases, one can check that if $Q$ contains no zero row/column, then $S\in\mathbb{S}_{+}^{n}$ and $J\in\mathbb{S}^{n}_{++}$, respectively.
Remark 3.1.
While our framework does not explicitly incorporate the same preconditioning techniques used in the conjugate gradient method, we argue that the matrix $J$ can be interpreted as a preconditioner. In this work, we primarily focus on block-diagonal matrices for $J$, as they enable efficient parallel implementation. However, more advanced strategies can be developed, provided that $J$ is invertible, and the matrix S=JQS=J-Q remains positive semidefinite. The design of $J$ is highly problem-specific and depends on the underlying structure and requirements of the application. As a consequence, exploring potential problem structures within a specific application domain can be a more effective approach.
@@ 4 Numerical experiments
In this section, we conduct numerical experiments to verify the efficiency and scalability of the proposed algorithms. The outline and main goals of our numerical experiments are summarized as follows:
@@
    We first construct a set of simple strictly diagonal dominant systems for which the classical Jacobi and weighted Jacobi methods are both guaranteed to be convergent. The purpose of this set of numerical tests is to show that our proposed algorithms outperform the classical Jacobi-type methods.
@@
    We repeat the above tests on some symmetric positive definite systems with real data sets in order to evaluate the performance of the proposed algorithm on symmetric positive definite systems.
@@
In our experiment, we only test Algorithm 3 with $J$ chosen as in (7), since, based on our numerical experience, Algorithm 3 with $J$ chosen as in (8) shares a similar performance as the former one. We terminate all the tested methods based on the value of the relative objective gap, i.e., $\frac{|f(x^{t})-f^{*}|}{1+|f^{*}|}$, where $x^{t}$ is the generated iterate at $t$-th iteration of an algorithm. In particular, we set the termination tolerance tol to $10^{-10}$. We also set the maximum number of iterations allowed, maxiter, to be $5000$.
The update rules for each algorithm are summarized in Table 1. For the accelerated Jacobi method, we choose $J$ to be a diagonal matrix. As demonstrated, the differences in per-iteration computational costs are negligible, making a comparison based solely on the number of iterations sufficient. Additionally, the practical performance heavily depends on the implementation of matrix-vector multiplications. Leveraging state-of-the-art implementations can further optimize performance, but this aspect lies beyond the scope of this paper. In view of the above, we have excluded a comparison of wall-clock time.
Method Update ($t\geq 0$) Complexity
Jacobi $x^{t+1}=x^{t}+D^{-1}(b-Qx^{t})$ $O(n^{2})$
Weighted Jacobi $x^{t+1}=x^{t}+\omega D^{-1}(b-Qx^{t})$ $O(n^{2})$
Accelerated Jacobi $x^{t+1}=y^{t}+J^{-1}(b-Qx^{t})$ $O(n^{2})$
$y^{t+1}=\left(1+\frac{\alpha_{t}-1}{\alpha_{t+1}}\right)x^{t}-\frac{\alpha_{t}-1}{\alpha_{t+1}}x^{t-1}$
Conjugate Gradient $x^{t+1}=x^{t}+\frac{\langle r^{t},r^{t}\rangle}{\langle p^{t},Qp^{t}\rangle}p^{t}$ $O(n^{2})$
$r^{t+1}=r^{t}-\frac{\langle r^{t},r^{t}\rangle}{\langle p^{t},Qp^{t}\rangle}Qp^{t}$
$p^{t+1}=r^{t+1}+\frac{\langle r^{t+1},r^{t+1}\rangle}{\langle r^{t},r^{t}\rangle}p^{t}$
Table 1: Testing methods and their per-iteration complexities.
@@ 4.1 Strictly diagonal dominant systems
Figure 2: Computational results for strictly diagonally dominant systems.
From Figure 2, we observe that both the Jacobi method and the weighted Jacobi method are convergent and admit linear convergence rates. Moreover, the weighted Jacobi method with the optimal weight has a faster convergence rate than that of the classical Jacobi method. However, both methods are not able to solve most problems to the desired accuracy within 5000 iterations. On the other hand, the proposed algorithms, i.e., Algorithm 1 and Algorithm 2 significantly outperform the other two methods and are able to solve all the problems successfully to the desired accuracy. Furthermore, by leveraging the adapative restarting scheme, the overall performance is improved significantly. These provide strong evidence that the Nesterov’s accelerated technique equipped with the adaptive restarting starting strategy is indeed efficient.
@@ 4.2 Laplacian systems
$L(G)]_{ij}:=\left\{\begin{array}[]{ll}\textrm{deg}(v_{i}),&\textrm{if $i=j$},\\ -1,&\textrm{if $i\neq j$ and $(i,j)\in E$},\\ 0,&\textrm{otherwise,}\end{array}\right.\quad 1\leq i,j\leq n,$
where $\textrm{deg}(v_{i})$ denotes the degree of the node $v_{i}$, for $1\leq i\leq n$. It is well-known that the Laplacian matrix $L(G)$ is symmetric positive semidefinite and singular with $\textrm{rank}(L(G))=n-1$. Solving Laplacian linear systems is a fundamental computational task with fruitful applications in a wide range of fields. The classical Jacobi method and the weighted Jacobi method are not guaranteed to be convergent when solving Laplacian systems. In fact, our numerical tests show that both methods are divergent empirically. Therefore, we only compare the proposed Algorithm 2 with the CG method and the PCG method with a simple diagonal preconditioner. We select a set of graphs arising from real-world applications in the SuiteSparse Matrix Collection  [^2] Available at https://sparse.tamu.edu/. with various sizes. The corresponding numerical results are presented in Figure 3.
@@
Figure 3: Computational results for Laplacian systems.
We can observe from Figure 3 that all the three methods are quite efficient and are able to solve all the tested problems to the specified accuracy. We can also see that Algorithm 2 and PCG have shown superior performance when compared to the pure CG method, i.e., CG. However, Algorithm 2 performs slightly worse than the PCG method. In particular, the number of iterations taken by Algorithm 2 is sometimes more than twice the number of iterations taken by PCG.
4.3 Symmetric positive definite systems
We also test Algorithm 2, CG and PCG on symmetric positive definite systems whose coefficient matrices are also collected from the SuiteSparse Matrix Collection. These matrices are real-world data coming from different practical science and engineering problems. The numerical results presented in Figure 4 show that Algorithm 2 performs worse than PCG and generally performs better than CG.
To assess the impact of the preconditioner, we also tested PCG equipped with the same diagonal preconditioner $J$ used in our proposed method, where $J$ is defined as the vector of row (or column) absolute sums of the matrix $Q$. Based on our observation, the resulting variant exhibits performance nearly identical to that of the vanilla CG method. For simplicity, the detailed computational results are not included in the paper. This suggests that the observed performance differences are not primarily due to the choice of preconditioner. While using the diagonal of $Q$ as $J$ may seem natural, it can violate the condition required by our analysis. However, problem-specific choices of $J$ may offer further improvements in efficiency.
(a) bcsstk13
(b) ex13
(c) msc01050
(d) msc01440
(e) msc10848
(f) nasa1824
(g) nasa2146
(h) nasa2910
(i) Trefethen-2000
Figure 4: Computational results for symmetric positive definite systems collected from the SuiteSparse collection.